# User-customizable Shared Control for Fine Teleoperation via Virtual Reality

### v1

@@ IV-D User-customizable Arbitration
Figure 4: The customizable spider chart interface in virtual reality. The operator can view and modify assistance across five factors, before confirming their configuration. Each factor is scaled from $[0.1,1]$ to $[10,100]$ with a discrete minimal update step of $5$.
A VR interface acts as the basis for bidirectional communication, $\bm{\psi}(\mathbf{v}_{r},v_{h})$, in our user-customizable shared control. We created a spider chart interface, depicted in Fig. 4, to augment the robot’s feedback, $\mathbf{v}_{r}$, with a variety of human-readable indicators, $\mathcal{V}$. Specifically, three factors affecting arbitration are displayed in VR: speed, depth assistance, and turnability. Speed regulates robot velocity along axes of motion relevant to the buzz wire game (e.g., horizontally along the wire). Depth assistance governs the margin between the end-effector and the wire along the depth axis, with the aim of assisting line-of-sight occlusions. Turnability adjusts the rotational speed of the end-effector. Users can thus view $\mathbf{v}_{r}$, as well as modify the parameters, $\bm{\theta}$, by supplying edits, $v_{h}$, per factor via increment/decrement buttons.
Following the formulation in Eq. 7, the arbitration vector, $\bm{\alpha}\in\mathbb{R}^{6}$, is expressed as a linear combination:
@@
$\text{where}\quad\mathbf{W}$ $=\begin{bmatrix}1&0&1&0&0.4&0.2\\ 0&1&0&0.2&0&0\\ 0&0&0&0.8&0.6&0.8\end{bmatrix}^{\mathsf{T}}.$ (23)
The coefficient matrix, $\mathbf{W}\in\mathbb{R}^{6\times 3}$, is structured in a manner where each editable facet, $v_{h}$, directly maps to a $\theta$ with a semantically similar effect on arbitration. The weighting of elements was chosen empirically to fit the task. For example, $v_{h}^{\mathrm{turn}}$ is masked by $\mathbf{W}$ to contribute to an $80\%$ change in arbitration on rotation about the z-axis, while $v_{h}^{\mathrm{speed}}$ contributes to the remaining $20\%$. Here, the term turnability intuitively represents more influence over robot rotation than translational speed. The weighting criteria for depth assistance uses the complement of $v_{h}^{\mathrm{depth}}$, as “assistance” implies more contribution from the autonomy.
In addition to task-specific parameters $\bm{\theta}$, we expose two task-independent parameters: safety and responsiveness. These parameters adjust $f_{\mathrm{max}}$ in Eq. 12 and $\mathbf{k}$ in Eq. 10, respectively, akin to a variable impedance controller. Note that each configurable parameter is actually constrained in the range $[0.1,1]$ to prevent a “full autonomy” configuration.
@@ V Experiment
To gauge the effectiveness of the proposed framework in enhancing teleoperation performance, we conducted a between-subjects study involving 12 subjects (3 female; aged 20-34). Subjects were asked to complete the buzz wire game by teleoperating a robotic arm in VR. All participants provided written consent prior to data collection.
@@ V-A Experiment Protocol
Participants were allocated into one of three groups, each designed to evaluate a different control strategy: (1) direct teleoperation (teleop), which is implemented by setting $\mathbf{A}_{\bm{\theta}}\,{=}\,\mathbf{I}$ in Eq. 9; (2) Shared control with a heuristics-based arbitration function (sc), which is described in Section V-B; and (3) our user-customizable shared control method (sc_user). Participants were unaware of the distinctions between each group. It is also worth noting that half of the 12 subjects reported no background in robotics. This equal distribution persisted across the control strategies.
Before commencing the experiment, participants in groups sc and sc_user were provided with identical written explanations of the five arbitration factors, described in Section IV-D. After each trial, these participants could review the updated arbitration factors for the subsequent trial through the VR interface (Fig. 4). Only sc_user subjects could edit these factors. All subjects were instructed to complete the game as safely and quickly as possible.
Each group consisted of four participants, and every subject carried out five consecutive trials of the buzz wire game in a single session using their assigned control strategy. This procedure was repeated for a total of four sessions. To evaluate the learning rates of operators over an extended period, a minimum of one workday separated each session, resulting in a two-week duration for the entire experiment. Additionally, the last session involved a distinct wire testbed, as portrayed in Fig. 5. This alteration aimed to assess whether any teleoperation skills gained are transferable to a related, yet different task. Subjects also answered post-session surveys on their user experience.
@@ V-B Heuristics-based Shared Control
We implemented a heuristics-based assist-as-needed controller as the reference method, sc, for our experiments. The arbitration parameters, $\bm{\theta}$, in this controller are iteratively updated according to the operator’s performance after each trial. We linearly update the parameters with a change rate akin to the multi-trial update function from [24]:
@@
where $t$ is the trial index and $\bm{\chi}_{t}\in\mathbb{R}^{3}$ is the change rate vector. $\bm{\chi}_{t}$ is updated using a non-linear function based on the two most recent trials:
@@
where $r_{d}^{i}$ is the $i^{\mathrm{th}}$ component of the desired error vector $\mathbf{r}_{d}$, ${\bar{r}^{i}_{t}}$ is the $i^{\mathrm{th}}$ component of the average error vector ${\bar{\mathbf{r}}_{t}}$ during trial $t$, and $\chi_{\mathrm{nom}}$ is the nominal change rate, which determines the maximum change rate within one update step.
@@
where $\overline{\mathbf{w}}_{a,t}\in\mathbb{R}^{6}$ is the average wrench from trial $t$, $\mathbf{W}^{\dagger}$ is the pseudo-inverse of the coefficient matrix in Eq. 19, and $\mathbf{r}_{d}$ is pre-calculated from an expert demonstration.
@@ V-C Evaluation Metrics
Performance evaluation for the buzz wire game primarily revolves around two measures: successful task completion time and number of collisions. A trial is deemed successful and collisions are counted when the participant teleoperates the robot’s end-effector to the game end without encountering a fatal failure (when collision forces exceed a threshold). Aside from task time and collisions, we also examine a user’s jerk in handling the VR controller (2nd derivative of input velocity commands) to capture “smoothness” [31]. Additionally, we explore the learning trajectories of users, i.e., each user’s change in “skill” metrics relative to their first session. Lastly, we investigate a user’s perceptions of the different control modes in terms of skill, success, and difficulty.
@@ VI Results
In the following, we report on our results and test our three hypotheses. When evaluating the learning trajectories of participants, we term the first session as the ‘baseline’, the combined second and third sessions as ‘training’, and the last session as ‘transfer’. Unless otherwise stated, mixed ANOVAs were performed to analyze the effects of control modes (between-subjects) and sessions (within-subjects) on the relevant evaluation metrics. The ‘baseline’ session was left out of the ANOVAs, leaving only the combined ‘training’ sessions to compare against the ‘transfer’ condition.
@@ VI-A Subjective Results
TABLE II: Subject responses to overall experiment questions (1 = “strongly disagree”, 5 = “strongly agree”).
@@
Subject ratings to post-session questions on skill (“How would you rate your skill?”), success (“How successful were you?”), and difficulty (“How hard did you have to work?”) are shown in Table I. For the ‘baseline’ session, sc subjects self-reported better ratings than the other two groups, especially in terms of perceived difficulty, with a mean difference of $1.7$ between sc and sc_user. There is a tendency toward more favorable responses for teleop and sc_user in the ‘training’ sessions, with teleop receiving the best ratings across all categories. Nevertheless, sc_user has the superior ‘transfer’ according to perceived skill ($7.3\pm 0.5$) and success ($7.0\pm 0.8$). The teleop users found the new wire testbed most difficult to complete ($7.0\pm 3.6$), with a mean score difference of $2.0$ from sc and $1.2$ from sc_user. Despite these patterns, no main or interaction effects were revealed from the mixed ANOVAs performed on each question category. Based on these tests, we cannot substantiate H1.
Survey questions on the entire experiment are summarized in Table II. Indeed, the last session is voted as marginally more difficult by teleop participants. Subjects in the sc_user group also viewed operating the robot as “quicker” to grasp than the other modes. Interestingly, sc_user participants considered the spider chart to be more “helpful” ($4.3\pm 1.0$) than the sc chart ($3.5\pm 1.0$), which had autonomously computed values and could not be edited. All sc_user individuals “found the VR interface easy to edit” ($5.0$, not in table).
@@ VII Discussion
A few notable findings emerged from the buzz wire experiment. First, while the subjective and absolute quantitative results were comparable for sc_user and teleop during ‘training’ sessions, sc_user operators outperformed the other two control modes in measures of relative improvement. For example, sc_user subjects had consistently less collisions and produced less jerky control inputs than their ‘baseline’ session, which were statistically significant differences over teleop and sc in the ‘transfer’ condition. Indeed, sc_user obtained markedly better results across all subjective and quantitative (absolute and relative) metrics in the new game. Altogether, these results reinforce the benefits of offering users an intuitive means of customizing shared control, particularly when there are variations in the task.
@@
Another key observation is the inferior performance of the sc policy. Comparing the distributions of arbitration parameters between sc_user and sc reveals higher variance in the sc_user group’s final selection of $\bm{\theta}$, as depicted in Fig. 8. The invariance in sc is due to the performance heuristic $\bar{\mathbf{r}}_{t}$ from Eq. 26 being solely dependent on the distance between the end-effector and wire. An alternative update routine based on $\bar{\mathbf{r}}_{t}$ could prioritize factors that sc_user participants gravitated toward, such as turnability and responsiveness. However, any heuristics-based method for automatically adapting $\bm{\theta}$ would still converge to a similar parameter distribution across users, without accounting for individual preferences. Shared control policies learnt from large demonstration datasets, e.g., using reinforcement learning [32], would also suffer from the lack of personalization to unseen users. Moreover, labeling arbitration parameters for a dataset requires regular human intervention and is thus difficult to simulate or acquire.
Lastly, we believe our formulation for user-customizable shared control can be applied to a broad range of teleoperation contexts. This formulation not only modeled arbitration factors that are specific to the task (depth assistance, speed, and turnability), but also incorporated task-independent factors, such as safety to tune the potential force magnitude and responsiveness to adjust the stiffness in admittance control. One could also extend the proposed shared control formulation to treat the dimensions of latent actions learnt from expert demonstrations as user-adjustable [11, 32, 33]. The main challenge here would be in communicating these dimensions through $\mathcal{V}$, and supplying interface actions, $v_{h}$, to modify their values. Each independent latent dimension may not represent an interpretable factor of variation, nor occupy a constrained space of values for easy editing via $v_{h}$. We defer investigation into this idea for future work.
@@ VIII Conclusions
In this paper, we introduced a mathematical shared control framework for users to customize the arbitration process. We presented an instantiation of this framework for a teleoperation task in $\mathrm{SE}(3)$ and conducted a longitudinal study spanning two weeks per subject to evaluate the proposed method. Our user-customizable shared control method enhanced teleoperation outcomes, including collision frequency and input jerk, while also accommodating the baseline skill levels of different users. Furthermore, operators transferred to generalizations in the teleoperation task better when they could edit the inner workings of the shared control. We believe these preliminary results hold promise for long-term adaptability to diverse variations in users and tasks. Future research directions will increase the participant sample size, explore a wider array of teleoperation tasks, and consider more flexible representations of the arbitration parameters for communication and user editing.

### v2

@@ IV-D User-customizable Arbitration
directional communication, $\bm{\psi}(\mathbf{v}_{r},v_{h})$. We created a spider chart interface, depicted in Fig. 4, to augment the robot’s feedback, $\mathbf{v}_{r}$, with a variety of human-readable indicators, $\mathcal{V}$. Before introducing these indicators, we emphasize that the spider chart is only visible when a user wishes to edit the shared control. Hence, no extra cognitive workload is induced during a game trial.
Figure 4: The customizable spider chart interface in virtual reality. The operator can modify any of the five factors in steps of $5$ before confirming their choice. Each factor is scaled from $[0.1,1]$ to $[10,100]$.
The $\mathcal{V}$ set consists of three arbitration factors displayed in VR: speed, depth assistance, and turnability. Speed regulates translational velocity along axes of motion relevant to the buzz wire game (e.g., horizontally along the wire). Depth assistance governs the margin between the end-effector and wire along the depth axis, with the aim of assisting line-of-sight occlusions. Turnability adjusts the rotational speed of the end-effector. Users can therefore examine the current parameters, $\bm{\theta}$, via the spider chart, as well as supply edits per factor, $v_{h}$, in increments/decrements (+/- buttons in Fig. 4). Although we defined these three factors for the buzz wire game, regulating motion in relevant DoF is widely used in shared control [9] (e.g., virtual fixture methods), and maintaining a clear line-of-sight is crucial for robot teleoperation in general [18]. We thus believe these arbitration factors can be easily adapted to many different robot teleoperation tasks.
To apply user edits $v_{h}$ to the arbitration, we express $\bm{\alpha}\in\mathbb{R}^{6}$ from Eq. 4 as a linear combination:
@@
$\text{where}\quad\mathbf{W}$ $=\begin{bmatrix}1&0&1&0&0.4&0.2\\ 0&1&0&0.2&0&0\\ 0&0&0&0.8&0.6&0.8\end{bmatrix}^{\mathsf{T}}.$
The coefficient matrix, $\mathbf{W}\in\mathbb{R}^{6\times 3}$, is structured in a manner where each editable facet, $v_{h}$, directly maps to a parameter with a semantically similar effect on arbitration. The weighting of elements was chosen empirically to fit the task. For example, $v_{h}^{\mathrm{turn}}$ is masked by $\mathbf{W}$ to contribute to an $80\%$ change in arbitration on rotation about the z-axis (yaw), while $v_{h}^{\mathrm{speed}}$ contributes to the remaining $20\%$. Here, the term turnability intuitively represents more influence over robot rotation than translational speed. The weighting criteria for depth assistance uses the complement of $v_{h}^{\mathrm{depth}}$, as “assistance” implies more contribution from the autonomy.
We also expose two task-independent factors through $\mathcal{V}$: safety and responsiveness. These factors adjust $f_{\mathrm{max}}$ in Eq. 9 and $\mathbf{k}$ in Eq. 7, respectively, akin to a variable impedance controller. Using safety to tune the force magnitude and responsiveness to adjust the stiffness are generalizable to any admittance control scheme. Each of the five configurable factors in $\mathcal{V}$ is constrained in the range $[0.1,1]$ to prevent a “full autonomy” configuration.
@@ V Experiment
To gauge the effectiveness of the proposed framework, we conducted a between-subjects study involving 12 subjects (3 female; aged 20-34). Subjects were asked to complete the buzz wire game by teleoperating a robotic arm in VR. All participants provided written consent prior to data collection.
@@ V-A Experiment Protocol
Participants were allocated into one of three groups, each designed to evaluate a different control strategy: (1) direct teleoperation (teleop), which is implemented by setting $\mathbf{A}_{\bm{\theta}}\,{=}\,\mathbf{I}$ in Eq. 6; (2) shared control with a heuristics-based arbitration function (sc), as outlined in Section V-B; and (3) our user-customizable shared control method (sc_user). Participants were unaware of the distinctions between each group. Subjects in sc and sc_user were provided with identical written explanations of the five factors from Section IV-D before the experiment. For instance, “Increases/Decreases in this parameter mean more/less robot assistance in avoiding obstacles” explained safety. It is also worth noting that an equal distribution of the subject pool reported no background in robotics, which was preserved across control strategies.
Each group consisted of four participants and every subject carried out five consecutive game trials in a single session using their assigned control strategy. This procedure was repeated for a total of four sessions. Subjects were also offered five minutes before their first session to become familiar with the VR setup while not playing the game. This included a brief overview of the spider chart shown in Fig. 4 for the sc and sc_user groups. To evaluate the learning rates of operators over an extended period, a minimum of one workday separated each session, resulting in an experiment spanning two weeks. Additionally, the last session involved a distinct wire testbed, as portrayed in Fig. 5. This alteration aimed to assess whether any teleoperation skills gained are transferable to a related, yet different task. Subjects also answered post-session surveys on their user experience.
After each trial, participants could spend a minute operating the arm without playing the game. For the teleop group, this post-trial time could be used as an opportunity to improve their understanding of direct robot teleoperation. Participants in sc could instead examine the automatically updated arbitration factors (see Section V-B) through the spider chart and test out the resulting control. Only the sc_user group could use this post-trial time to edit the arbitration factors displayed on the spider chart and explore different configurations. User-selected arbitration settings persisted across trials unless altered. During game trials, the spider chart was not visible and arbitration settings were fixed.
@@ V-B Heuristics-based Shared Control
We implemented a heuristics-based assist-as-needed controller as the reference method, sc, for our experiment. The arbitration parameters, $\bm{\theta}$, in this controller are iteratively updated according to the operator’s performance after each trial. We linearly updated the parameters with a change rate akin to the multi-trial update function from [26]:
@@
where $t$ is the trial index and $\bm{\chi}_{t}\in\mathbb{R}^{3}$ is the change rate vector, which is updated using the two most recent trials:
@@
where $r_{d}^{i}$ is the $i^{\mathrm{th}}$ component of the desired error vector $\mathbf{r}_{d}$ (pre-calculated from an expert demonstration), ${\bar{r}^{i}_{t}}$ is the $i^{\mathrm{th}}$ component of the average error vector ${\bar{\mathbf{r}}_{t}}$ during trial $t$, and $\chi_{\mathrm{nom}}$ is the nominal change rate, which determines the maximum change rate within one update step.
@@
where $\overline{\mathbf{w}}_{a,t}\in\mathbb{R}^{6}$ is the average wrench from trial $t$ and $\mathbf{W}^{\dagger}$ is the pseudo-inverse of the coefficient matrix in Eq. 13.
@@ V-C Evaluation Metrics
Performance evaluation for the buzz wire game primarily revolves around two measures: successful task completion time and number of collisions. A trial is deemed successful and collisions are counted when the participant teleoperates the robot’s end-effector to the game end (a pre-defined location along the wire) without encountering a fatal failure, i.e., when collision forces exceed a threshold. Aside from task time and collisions, we also examine a user’s jerk (the mean squared measure from [32]) in handling the VR controller to capture smoothness or fluency. Moreover, we explore the learning trajectories of users, i.e., each user’s change in “skill” metrics relative to their first session. Lastly, we investigate a user’s perceptions of the different control modes in terms of skill, success, and difficulty.
@@ VI Results
In the following, we report on our results and test our hypotheses. When evaluating the learning trajectories of participants, we term the first session as the ‘baseline’, the combined second and third sessions as ‘training’, and the last session as ‘transfer’. Unless otherwise stated, mixed ANOVAs were performed to analyze the effects of control modes (between-subjects) and sessions (within-subjects) on the relevant evaluation metrics. The ‘baseline’ session was left out of the ANOVAs, leaving only the combined ‘training’ sessions to compare against the ‘transfer’ condition.
@@ VI-A Subjective Results
TABLE II: Subject responses to overall experiment questions (1 = “strongly disagree”, 5 = “strongly agree”)
@@
Subject ratings to post-session questions on skill (“How would you rate your skill?”), success (“How successful were you?”), and difficulty (“How hard did you have to work?”) are shown in Table I. For the ‘baseline’ session, sc subjects reported better ratings than the other groups, especially in terms of difficulty, with a mean difference of $1.7$ compared to sc_user. There is a tendency toward more favorable responses for teleop and sc_user in the ‘training’ sessions, with teleop receiving the best ratings across all categories. Nevertheless, sc_user has the superior ‘transfer’ according to perceived skill ($7.3\pm 0.5$) and success ($7.0\pm 0.8$). The teleop users found the new wire testbed more difficult to complete ($7.0\pm 3.6$), with a mean difference of $2.0$ from sc and $1.2$ from sc_user. Despite these patterns, no main or interaction effects were revealed from mixed ANOVAs performed on each question category and so we cannot substantiate H1.
Survey questions on the entire experiment are summarized in Table II. Indeed, the last session is voted as marginally more difficult by teleop participants. Subjects in the sc_user group also viewed operating the robot as quicker to learn than sc and even teleop. Interestingly, sc_user participants considered the spider chart to be more helpful ($4.3\pm 1.0$) than the sc chart ($3.5\pm 1.0$), which had autonomously computed values and could not be edited. All sc_user individuals “found the VR interface easy to edit” ($5.0$, not in table).
@@ VII Discussion
A few notable findings emerged from the buzz wire experiment. First, while the subjective and absolute results were comparable for sc_user and teleop during ‘training’ sessions, sc_user operators outperformed the other control modes in measures of relative improvement. For example, sc_user subjects had consistently less collisions and produced less jerky inputs than their ‘baseline’ session, which were statistically significant differences over teleop and sc in the ‘transfer’ condition. Indeed, sc_user obtained markedly better results across all subjective and quantitative (absolute and relative) metrics in the new game. Altogether, these results reinforce the benefits of offering users a means of customizing shared control, particularly when there are variations in the task.
@@
Another observation is the inferior performance of the sc policy. Comparing the distributions of arbitration parameters between sc_user and sc reveals higher variance in the sc_user group’s final $\bm{\theta}$ selection, as depicted in Fig. 8. The invariance in sc is due to the performance heuristic $\bar{\mathbf{r}}_{t}$ from Eq. 16 being solely dependent on the distance between the end-effector and wire. An alternative update routine could prioritize factors that sc_user participants gravitated toward, such as turnability and responsiveness. However, any heuristics-based method for automatically adapting $\bm{\theta}$ would still converge to a similar parameter distribution across users, without accounting for individual preferences. Shared control policies learned from large demonstration datasets would also suffer from the lack of personalization to unseen users.
Despite our promising preliminary findings, there are areas for improvement. One important limitation is the broader applicability of our results due to the constrained task and environment. Future research should explore more complex telemanipulation tasks, e.g., those involving humanoids [1], and include a larger, more diverse participant pool with varied backgrounds and skill levels in robot teleoperation. Another imperative requirement going forward will be to ensure that users have a clear understanding of the implications in visualization signals, $\mathcal{V}$, and interface actions, $v_{h}$. Regardless of our best efforts to make $\mathcal{V}$ and $v_{h}$ intuitive, there was one occasion where an sc_user subject misunderstood the meaning of a parameter, leading to a deterioration in performance. Simplifying the selection of arbitration factors may alleviate this issue, especially since certain factors, e.g., turnability, were redundant as all sc_user subjects converged on a similar setting (see Fig. 8).
Finally, we suggest ways to extend our user-customizable formulation to make it more widely applicable. Instead of using various task-specific factors, like speed, a single scalar $\alpha\in[0,1]$ for “autonomous assistance” could be easily tuned in VR, e.g., with a graphical slider. Modern frameworks for shared control, such as those based on reinforcement learning [33], could then let users calibrate $\alpha$ online, rather than relying on a fixed or learned value that they cannot see or refine. Alternatively, latent actions learned from expert demonstrations could be made user-adjustable [13, 34]. The main challenges here would be in communicating the latent dimensions through $\mathcal{V}$ and exposing user interface actions to modify their values, as these latent actions may not represent a human-interpretable factor or be easy to edit via $v_{h}$. We defer investigation into these ideas for future work.
@@ VIII Conclusions
In this paper, we introduced a mathematical shared control framework for users to customize the arbitration process. We presented an instantiation of this framework for a teleoperation task in $\mathrm{SE}(3)$ and conducted a longitudinal study spanning two weeks per subject to evaluate the proposed method. Our user-customizable shared control method enhanced teleoperation outcomes, reducing collision frequency and input jerk, while also accommodating the baseline skill levels of different users. Furthermore, operators transferred to generalizations in the teleoperation task better when they could edit the inner workings of the shared control. We believe these results hold promise for long-term adaptability to diverse variations in users and tasks. Future research will extend the scale of the user study, explore a wider array of tasks, and consider more flexible representations of the arbitration parameters for communication and user editing.